# Undefinability of Approximation of 2-to-2 Games

### v1

@@ 4.4 Shuffling variables
One issue that arises with the games constructed in the reduction from Section 3 is that we have a fixed alphabet of size $q=2^{l}$ and we associate with each vertex $(U,L)$ an arbitrary bijection between this and the $2^{l}$ distinct linear functions on the space $L+H_{U}$ that satisfy the equations in $U$. The consistency across different vertices is then enforced by the constraint relations. In order to turn this into a first-order reduction, we want to choose these bijections in a symmetry-preserving fashion.
Let $I$ be our starting instance of 3XOR and $I_{2:2}^{T}=\Theta(I)$ be the transitive $2$-to-$2$ game obtained from the first step of the reduction of Section 3, and let $X$ be the set of variables of $I$. Let $\rho\in\text{Sym}_{X}$ be a permutation of $X$. This permutation has a natural action on other objects constructed from $X$. In particular, for an equation $e$ of the form $x+y+z=b$, we write $\rho(e)$ for the equation $\rho(x)+\rho(y)+\rho(z)=b$. When $U$ is a tuple of such equations, we write $\rho(U)$ for the tuple obtained by applying $\rho$ componentwise to each element of the tuple. Similarly, for other objects obtained by set and tuple constructions from $X$, we apply the permutation $\rho$ to denote the natural induced action without defining it formally.
@@
Now, we want to choose the bijections between our set of $2^{l}$ labels and the linear functions associated with a vertex $(U,L)$ in such a way that whenever $(U,L)$ and $\rho(U,L)$ are both vertices in $I_{2:2}^{T}$, then they commute with $\rho$. For this, fix a canonical space $\mathbb{F}_{2}^{3k}$ of dimension $3k$. For each $U\in\mathcal{U}$, we write $X_{U}\subseteq X$ for the set of variables that appear in $U$. Since $U$ is a sequence of $k$ equations with pairwise disjoint sets of variables, we can fix a bijection between $X_{U}$ and $[3k]$ which induces an isomorphism $\mu_{U}:\mathbb{F}_{2}^{X_{U}}\rightarrow\mathbb{F}_{2}^{3k}$. These bijections can be chosen to be $\rho$-invariant (for all $\rho$), that is,
@@ 4.5 The reduction to the transitive game
We now describe how the reduction $\Theta$ from Section 3.2 can be given as a first-order interpretation. Fix positive integers $k$ and $l$, which are the parameters to the reduction. Given a (regular) 3XOR instance $\mathbb{A}=(X,\mathrm{Eq}_{0}^{\mathbb{A}},\mathrm{Eq}_{1}^{\mathbb{A}})$, our interpretation maps it to the following (transitive) $2$-to-$2$ game (with alphabet size $2^{l}$) $\mathbb{B}$.
@@
  • The next $k$ elements $r_{1},\dots,r_{k}$ define the right-hand sides of the $k$ equations in $U$. To encode these as binary values, we use $r_{i}=u_{1,1}$ to encode the value 00 and $r_{i}=u_{1,2}$ to encode the value $1$. Since $u_{1,1}$ and $u_{1,2}$ are distinct, this works and can be specified by a first-order formula.
  • The next $2^{3k}$ elements also encode bits, using the values of $u_{1,1}$ and $u_{1,2}$ as 00 and $1$. Think of these as specifying a subset of $\mathbb{F}_{2}^{3k}$. We can write a first-order formula that says that this subset is a subspace $L^{*}$ of dimension $l$ (since $l$ and $k$ are fixed, the formula is simply a big disjunction over all subspaces). Finally, we can also write a first-order formula that checks that $L^{*}$ is in $\mathcal{L}$.
@@
Given two vertices $(U,L)$ and $(U^{\prime},L^{\prime})$ of $\mathbb{B}$, the type of constraint between them ($1$-to-$1$, $2$-to-$2$ or no constraint at all) only depends on $\mu_{U}(L),\mu_{U^{\prime}}(L^{\prime}),r,r^{\prime}$ and $I(U,U^{\prime})$, where $r$,$r^{\prime}$ are the vectors of the right-hand sides of the equations and
@@ 4.6 Weight approximation
The vertices of the integer-weighted game are exactly those in the structure $\mathbb{B}$ above. The main task is to define the weights, by defining a suitable set $C$ of constraints. Recall that the vertices of $I^{w}_{2:2}$ are partitioned into cliques $C_{1},\dots,C_{m}$ based on the $1$-to-$1$ constraints. Suppose $(U_{1},L_{1})\in C_{i}$ and $(U_{2},L_{2})\in C_{j}$ are two vertices connected by a $2$-to-$2$ constraint. Then, the weight of the constraint is
@@
Given a weighted 2-to-2 game $G=(V,\Sigma,\Phi,w)$, whose value is at most $\delta$, any game $G^{\prime}=(V,\Sigma,\Phi,w^{\prime})$ where $\forall\phi\in\Phi.\frac{1}{\gamma}<\frac{w(\phi)}{w^{\prime}(\phi)}<\gamma$ has value at most $\delta\gamma^{2}$.
@@
(sketch) The sum of weights drops at most by a factor $\gamma$, and the sum of the weights of the satisfied constraints increases by at most a factor of $\gamma$. ∎
@@
The number of elements in an equivalence class with $f$ useful equations is simply the number of ways to set the remaining $k-f$ equations. This can be approximated by $|\mathrm{Eq}|^{k-f}$. Given $f$ useful equations, the probability of a random set of $k-f$ equations having common variables with $U$, the set of useful equations or each other, or making the $k$-tuple invalid by having two variables from different equations which have a common equation in the 3XOR instance, converges to zero ($O\left(\frac{k^{2}}{|X|}\right)$) as the instance size grows, due to the regularity condition. By adding all the approximate sizes of the equivalence classes within $C_{i}$, we can conclude that the approximation
@@
$\begin{split}w((U_{1},&L_{1}),(U_{2},L_{2}))=|\{(U,L,L^{\prime})\mid(U,L)\in C_{i},(U,L^{\prime})\in C_{j}\}|\\ &\cdot\prod_{\bm{v}\in\{0,\dots,\Psi\}^{k+1}}\begin{cases}\chi(\bm{v})&\text{if }\bm{v}\neq\bm{\nu_{(U_{1},L_{1})}}\\ 1&\text{if }\bm{v}=\bm{\nu_{(U_{1},L_{1})}}\end{cases}\cdot\prod_{\bm{v}\in\{0,\dots,\Psi\}^{k+1}}\begin{cases}\chi(\bm{v})&\text{if }\bm{v}\neq\bm{\nu_{(U_{2},L_{2})}}\\ 1&\text{if }\bm{v}=\bm{\nu_{(U_{2},L_{2})}}\\ \end{cases}\end{split}$ (2)
As we see next, we can define a reduction in $\mathrm{FO}$ to $2$-to-$2$ games using these approximate weights.
@@ 4.7 Defining the unweighted game
Finally, we are ready to show that the construction of an unweighted edge consistent $2$-to-$2$ game (with multiple edges to approximate weights as above) can be given by an $\mathrm{FO}$ interpretation.
@@
A vertex $(U,L)$ is coded by the first $4k+2^{3k}$ elements of this tuple, as before, followed by a sequence of 00s. Recall that we code bits 00 and $1$ by the first and second elements of the tuple. The first of these 00s is to be interpreted as an indicator that the tuple is a vertex (it will be $1$ for a constraint), and the rest are padding to make the length of the tuples match.
A constraint $c$ is coded by a tuple where the first $4k+2^{3k}$ elements represent a vertex $(U,L)$, this is followed by a $1$ (i.e. a repeat of the second element of the tuple) and then the next $4k+2^{3k}$ represent a second vertex $(U^{\prime},L^{\prime})$. The rest of the tuple codes a unique identifier of the constraint, $\mathrm{ID}$. We construct the interpretation so that for all fixed $(U,L),(U^{\prime},L^{\prime})$, there are $w((U,L),(U^{\prime},L^{\prime}))$ different identifiers where $w$ is the approximate weight described above. We show that for this weight function, there is a formula $W$ which defines a set of exactly $w((U,L),(U^{\prime},L^{\prime}))$ tuples extending the description of $(U,L)$ and $(U^{\prime},L^{\prime})$.
@@
Thus, we can define the formulas defining the set of vertices and constraints. For simplicity, we use $U,L,U^{\prime},L^{\prime},\mathrm{ID}$ to describe the sub-tuple of variables in their corresponding parts of the $N$-tuple, where $N=8k+1+2^{3k+1}+Q$.
$\mathrm{Node}(U,L,\mathrm{IsConstraint},U^{\prime},L^{\prime},\mathrm{ID})\equiv$ $\mathrm{IsConstraint}=0\wedge\pi^{U}{(U,L)}\wedge\operatorname*{\wedge}_{x\in(U^{\prime},L^{\prime},\mathrm{ID})}x=0$
To check if it is a valid constraint, we need
$\mathrm{Constraint}(U,L,\mathrm{IsConstraint},U^{\prime},L^{\prime}$ $,\mathrm{ID})\equiv\mathrm{IsConstraint}=1\wedge\pi^{U}(U,L)\wedge\pi^{U}(U^{\prime},L^{\prime})$
$\wedge\operatorname*{\vee}_{C\in\mathcal{C}_{2}}\pi^{C}((U,L),(U^{\prime},L^{\prime}))\wedge W((U,L),(U^{\prime},L^{\prime}),\mathrm{ID})$
@@ 4.8 Defining W
To define $W$, we construct formulas defining various elements of Equation 2. More precisely, for various numerical expressions $e(x,y)$, which depend on the values assigned to $x$ and $y$, we construct formulas we denote $w_{q,e}(x,y,z)$, where $q$ is the length of the tuple of variables $z$. These formulas have the property that when $x$ and $y$ are interepreted by the elements coding the nodes $(U,L)$ and $(U^{\prime},L^{\prime})$ the number of $q$-tuples that can be assigned to $z$ to make $\omega_{q,e}$ true is exactly $e(x,y)$. As before, we use 00 and $1$ to denote the first and second elements of the tuple. Also, for a first-order formula $\phi(x,y)$, let $1_{\phi}$ denote the indicator variable that $\phi$ is true (under an assignment of values to $x$ and $y$).
$e=1$: $\omega_{1,e}(x,y,z)\equiv(z=0)$
$e=1_{\phi}$: $\omega_{1,e}(x,y,z)\equiv(z=0)\wedge\phi(x,y)$
@@
$\omega_{q_{1}+q_{2},e}(x,y,z_{1},\dots,z_{q_{1}},z_{q_{1}+1},\dots,z_{q_{2}})\equiv\omega_{q_{1},e_{1}}(z_{1},\dots,z_{q_{1}})\wedge\omega_{q_{2},e_{2}}(z_{q_{1}+1},\dots,z_{q_{2}})$
@@
$\bigvee\left[z_{1}=1\bigwedge\omega_{q_{2},e_{2}}(z_{2},\dots,z_{q_{2}+1})\right]$
@@
$\omega_{4,e}(x,y,z_{1},z_{2},z_{3},z_{4})\equiv\left(z_{1}=0\land\mathrm{Eq}_{0}(z_{2},z_{3},z_{4})\right)\vee\left(z_{1}=1\land\mathrm{Eq}_{1}(z_{2},z_{3},z_{4})\right)$
@@
$\wedge\operatorname*{\bigvee}_{C\in\mathcal{C}_{1}}C((z_{1},z_{2}),x)\wedge\operatorname*{\bigvee}_{C\in\mathcal{C}_{1}}C((z_{1},z_{3}),y)$
@@ 5.2 Vertex Cover
To prove Theorem 5.2, it suffices to show that the reduction in [24, Chapter 5] from $\mathrm{Gap}\textsc{Irreg2to2}_{j,q}(1-\epsilon,\delta^{\prime})$ to $\mathrm{GapIS}(1-\frac{1}{\sqrt{2}}-\delta,\epsilon)$ is definable in $\mathrm{FO}$ for any $\epsilon,\delta$ with a suitable choice of $j,q,\delta^{\prime}$ such that $\mathrm{Gap}\textsc{Irreg2to2}_{j,q}(1-\epsilon,\delta^{\prime})$ is undefinable (by Theorem 3.5). We first summarise the reduction and then show its $\mathrm{FO}$-definability. Note that, though Theorem 3.5 is proved with perfect completeness, this does not yield perfect completeness in Theorem 5.2, since the slack $\delta$ is introduced in the reduction.
We now give a description of the reduction that takes a $2$-to-$2$-game instance $((V,E),\Sigma,\Phi)$ to a node-weighted graph $(V^{\prime},E^{\prime},w:V^{\prime}\rightarrow\mathbb{Q}^{+})$ as follows.
@@
$E^{\prime}=\{((u,A_{1}),(v,A_{2})\mid(A_{1}\times A_{2})\cap\Phi(e)=\emptyset\text{ for an edge }e\text{ on }u,v\}$
We can understand this as giving us a graph whose vertices are nodes of the game, labelled by a set of possible colourings. The edges ensure that two vertices cannot be in an independent set together if no possbile colourings of the game nodes from the associated set satisfies the corresponding constraint. As shown in [24, Chapter 5], this is a reduction from $\mathrm{Gap}\textsc{Irreg2to2}_{j,q}(1-\epsilon^{\prime},\delta^{\prime})$ to $\mathrm{GapIS}(1-\frac{1}{\sqrt{2}}-\delta,\epsilon)$.
@@
We now show that this reduction can be defined as a first-order interpretation. The dimension of the reduction is $d=3+|\Sigma|+Q^{|\Sigma|}$. A tuple $t\in V^{d}$ is to be understood as coding one instance of the pair $(v,A)$. As before, the first two elements of $t$ are used to code bits 00 and $1$. The third component is the vertex $v$. The next $|\Sigma|$ components are bits coding the set $A$, and the final $Q^{|\Sigma|}$ elements are a binary identifier giving a value up to $W(v,A)$, thus ensuring that we have $W(v,A)$ instances of the vertex $(v,A)$.
@@
To define the edge relation, consider any $2$-to-$2$ relation $\Pi\subseteq\Sigma\times\Sigma$. We construct a first-order formula $\Pi(x,y)$, where $x$ and $y$ are $d$-tuples, that says that if $x$ and $y$ code vertices $(u,A)$ and $(v,B)$ then there are elements $s\in A$ and $t\in B$ such that $(s,t)\in\Pi$.
@@
$\phi_{E}(x,y):=\exists zC(z)\land\Phi_{\Pi}(x_{3},y_{3},z)\land\Pi(x,y).$

### v2

@@ 4.4 Shuffling variables
One issue that arises with the games constructed in the reduction from Section 3 is that we have a fixed alphabet of size $q=2^{l}$ and we associate with each vertex $(U,L)$ an arbitrary bijection between the alphabet and the $2^{l}$ distinct linear functions on the space $L+H_{U}$ that satisfy the equations in $U$. The consistency across different vertices is then enforced by the constraint relations. In order to turn this into a first-order reduction, we want to choose these bijections in a symmetry-preserving fashion.
Let $I$ be our starting instance of 3XOR and $I_{2:2}^{T}=\Theta(I)$ be the transitive 22-to-22 game obtained from the first step of the reduction of Section 3, and let $X$ be the set of variables of $I$. Let $\rho\in\text{Sym}_{X}$ be a permutation of $X$. This permutation has a natural action on other objects constructed from $X$. In particular, for an equation $e$ of the form $x+y+z=b$, we write $\rho(e)$ for the equation $\rho(x)+\rho(y)+\rho(z)=b$. When $U$ is a tuple of such equations, we write $\rho(U)$ for the tuple obtained by applying $\rho$ componentwise to each element of the tuple. Similarly, for other objects obtained by set and tuple constructions from $X$, we apply the permutation $\rho$ to denote the natural induced action without defining it formally.
@@
Now, we want to choose the bijections between our set of $2^{l}$ labels and the linear functions associated with a vertex $(U,L)$ in such a way that whenever $(U,L)$ and $\rho(U,L)$ are both vertices in $I_{2:2}^{T}$, then they commute with $\rho$. For this, fix a canonical space $\mathbb{F}_{2}^{3k}$ of dimension $3k$. For each $U\in\mathcal{U}$, we write $X_{U}\subseteq X$ for the set of variables that appear in $U$. Since $U$ is a sequence of $k$ equations with pairwise disjoint sets of variables, we can fix a bijection between $X_{U}$ and $[3k]$ which induces an isomorphism $\mu_{U}:\mathbb{F}_{2}^{X_{U}}\rightarrow\mathbb{F}_{2}^{3k}$. We can naturally extend $\mu_{U}$ to a map between subspaces: for a subspace $L\subseteq\mathbb{F}_{2}^{X_{U}}$, $\mu_{U}(L)=\{\mu_{U}(v)\mid v\in L\}$. These bijections can be chosen to be $\rho$-invariant (for all $\rho$), that is,
@@ 4.5 The reduction to the transitive game
We now describe how the reduction $\Theta$ from Section 3.2 can be given as a first-order interpretation. Fix positive integers $k$ and $l$, which are the parameters to the reduction. Given a (regular) 3XOR instance $\mathbb{A}=(X,\mathrm{Eq}_{0}^{\mathbb{A}},\mathrm{Eq}_{1}^{\mathbb{A}})$, our interpretation maps it to the following (transitive) 22-to-22 game (with alphabet size $2^{l}$) $\mathbb{B}$.
@@
  • The next $k$ elements $r_{1},\dots,r_{k}$ define the right-hand sides of the $k$ equations in $U$. To encode these as binary values, we use $r_{i}=u_{1,1}$ to encode the value 0 and $r_{i}=u_{1,2}$ to encode the value $1$. Since $u_{1,1}$ and $u_{1,2}$ are distinct, this works and can be specified by a first-order formula.
  • The next $2^{3k}$ elements also encode bits, using the values of $u_{1,1}$ and $u_{1,2}$ as 0 and $1$. Think of these as specifying a subset of $\mathbb{F}_{2}^{3k}$. We can write a first-order formula that says that this subset is a subspace $L^{*}$ of dimension $l$ (since $l$ and $k$ are fixed, the formula is simply a big disjunction over all subspaces). Finally, we can also write a first-order formula that checks that $L^{*}$ is in $\mathcal{L}$.
@@
Given two vertices $(U,L)$ and $(U^{\prime},L^{\prime})$ of $\mathbb{B}$, the type of constraint between them (11-to-11, 22-to-22 or no constraint at all) only depends on $\mu_{U}(L),\mu_{U^{\prime}}(L^{\prime}),r,r^{\prime}$ and $I(U,U^{\prime})$, where $rr^{\prime}$ are the vectors of the right-hand sides of the equations and
@@ 4.6 Weight approximation
The vertices of the integer-weighted game are exactly those in the structure $\mathbb{B}$ above. The main task is to define the weights, by defining a suitable set $C$ of constraints. Recall that the vertices of $I^{w}_{2:2}$ are partitioned into cliques $C_{1},\dots,C_{m}$ based on the 11-to-11 constraints. Suppose $(U_{1},L_{1})\in C_{i}$ and $(U_{2},L_{2})\in C_{j}$ are two vertices connected by a 22-to-22 constraint. Then, the weight of the constraint is
@@
Given a weighted 2-to-2 game $(G,\Sigma,\Phi,w)$, whose value is at most $\delta$, any game $(G,\Sigma,\Phi,w^{\prime})$ where $\forall e\in E(G).\frac{1}{\gamma}<\frac{w(e)}{w^{\prime}(e)}<\gamma$ has value at most $\delta\gamma^{2}$.
@@
The total sum of weights $\sum_{e\in E(G)}w^{\prime}(e)\geq\frac{1}{\gamma}\sum_{e\in E(G)}w(e)$ since $\frac{w(e)}{w^{\prime}(e)}<\gamma$ for all $e$. On the other hand, for any colouring $\chi:V(G)\rightarrow\Sigma$, we have $\sum_{e\in S_{\chi}}w^{\prime}(e)\leq\gamma\sum_{e\in S_{\chi}}w(e)$ and the result follows. ∎
@@
The number of elements in an equivalence class with $f$ useful equations is simply the number of ways to set the remaining kfk-f equations. This can be approximated by $|\mathrm{Eq}|^{k-f}$. Given $f$ useful equations, the probability of a random set of kfk-f equations having common variables with $U$, the set of useful equations or each other, or making the $k$-tuple invalid by having two variables from different equations which have a common equation in the 3XOR instance, converges to zero ($O\left(\frac{k^{2}}{|X|}\right)$) as the instance size grows, due to the regularity condition. By adding all the approximate sizes of the equivalence classes within $C_{i}$, we can conclude that the approximation
@@
[…]
As we see next, we can define a reduction in $\mathrm{FO}$ to 22-to-22 games using these approximate weights.
@@ 4.7 Defining the unweighted game
Finally, we are ready to show that the construction of an unweighted edge consistent 22-to-22 game (with multiple edges to approximate weights as above) can be given by an $\mathrm{FO}$ interpretation.
@@
A vertex $(U,L)$ is coded by the first $4k+2^{3k}$ elements of this tuple, as before, followed by a sequence of 0s. Recall that we code bits 0 and $1$ by the first and second elements of the tuple. The first of these 0s is to be interpreted as an indicator that the tuple is a vertex (it will be $1$ for a constraint), and the rest are padding to make the length of the tuples match.
A constraint $c$ is coded by a tuple where the first $4k+2^{3k}$ elements represent a vertex $(U,L)$, this is followed by a $1$ (i.e. a repeat of the second element of the tuple) and then the next $4k+2^{3k}$ elements represent a second vertex $(U^{\prime},L^{\prime})$. The rest of the tuple codes a unique identifier of the constraint, $\mathrm{ID}$. We construct the interpretation so that for all fixed $(U,L),(U^{\prime},L^{\prime})$, there are $w((U,L),(U^{\prime},L^{\prime}))$ different identifiers where $w$ is the approximate weight described above. We show that for this weight function, there is a formula $W$ which defines a set of exactly $w((U,L),(U^{\prime},L^{\prime}))$ tuples extending the description of $(U,L)$ and $(U^{\prime},L^{\prime})$.
@@
Thus, we can define the first-order formulas $\mathrm{Node}$ and $\mathrm{Constraint}$ defining the set of vertices and constraints. For simplicity, we use $U,L,U^{\prime},L^{\prime},\mathrm{ID}$ to describe the sub-tuple of variables in their corresponding parts of the $N$-tuple, where $N=8k+1+2^{3k+1}+Q$. Thus, the following formula defines exactly those tuples that are nodes.
$\mathrm{Node}(U,L,\mathrm{IsConstraint},U^{\prime},L^{\prime},\mathrm{ID})\equiv\mathrm{IsConstraint}=0\wedge\pi^{U}{(U,L)}\wedge\bigwedge_{x\in(U^{\prime},L^{\prime},\mathrm{ID})}x=0.$
The next formula defines exactly those tuples that are constraints.
$\mathrm{Constraint}$ $(U,L,\mathrm{IsConstraint},U^{\prime},L^{\prime},\mathrm{ID})\equiv$
$\mathrm{IsConstraint}=1\wedge\pi^{U}(U,L)\wedge\pi^{U}(U^{\prime},L^{\prime})\wedge\bigvee_{C\in\mathcal{C}_{2}}\pi^{C}((U,L),(U^{\prime},L^{\prime}))\wedge W((U,L),(U^{\prime},L^{\prime}),\mathrm{ID}).$
@@ 4.8 Defining W
To define $W$, we construct formulas defining various elements of Equation 2. More precisely, for various numerical expressions $e(x,y)$, which depend on the values assigned to $x$ and $y$, we construct formulas we denote $w_{q,e}(x,y,z)$, where $q$ is the length of the tuple of variables $z$. These formulas have the property that when $x$ and $y$ are interepreted by the elements coding the nodes $(U,L)$ and $(U^{\prime},L^{\prime})$ the number of $q$-tuples that can be assigned to $z$ to make $\omega_{q,e}$ true is exactly $e(x,y)$. As before, we use 0 and $1$ to denote the first and second elements of the tuple. Also, for a first-order formula $\phi(x,y)$, let $1_{\phi}$ denote the indicator variable that $\phi$ is true (under an assignment of values to $x$ and $y$).
$e=1$: $\omega_{1,e}(x,y,z)\equiv(z=0).$
$e=1_{\phi}$: $\omega_{1,e}(x,y,z)\equiv(z=0)\wedge\phi(x,y).$
@@
$\omega_{q_{1}+q_{2},e}(x,y,z_{1},\dots,z_{q_{1}},z_{q_{1}+1},\dots,z_{q_{2}})\equiv\omega_{q_{1},e_{1}}(z_{1},\dots,z_{q_{1}})\wedge\omega_{q_{2},e_{2}}(z_{q_{1}+1},\dots,z_{q_{2}}).$
@@
$\bigvee\left[z_{1}=1\bigwedge\omega_{q_{2},e_{2}}(z_{2},\dots,z_{q_{2}+1}).\right]$
@@
$\omega_{4,e}(x,y,z_{1},z_{2},z_{3},z_{4})\equiv\left(z_{1}=0\land\mathrm{Eq}_{0}(z_{2},z_{3},z_{4})\right)\vee\left(z_{1}=1\land\mathrm{Eq}_{1}(z_{2},z_{3},z_{4})\right).$
@@
$\wedge\operatorname*{\bigvee}_{C\in\mathcal{C}_{1}}C((z_{1},z_{2}),x)\wedge\operatorname*{\bigvee}_{C\in\mathcal{C}_{1}}C((z_{1},z_{3}),y).$
@@ 5.2 Vertex Cover
To prove Theorem 5.2, it suffices to show that the reduction in [26, Chapter 5] from
$\mathrm{Gap}\textsc{Irreg2to2}_{j,q}(1-\epsilon,\delta^{\prime})$ to $\mathrm{GapIS}(1-\frac{1}{\sqrt{2}}-\delta,\epsilon)$ is definable in $\mathrm{FO}$ for any $\epsilon,\delta$ with a suitable choice of $j,q,\delta^{\prime}$ such that $\mathrm{Gap}\textsc{Irreg2to2}_{j,q}(1-\epsilon,\delta^{\prime})$ is undefinable (by Theorem 3.5). We first summarise the reduction and then show its $\mathrm{FO}$-definability. Note that, though Theorem 3.5 is proved with perfect completeness, this does not yield perfect completeness in Theorem 5.2, since the slack $\delta$ is introduced in the reduction.
We now give a description of the reduction that takes a 22-to-22-game instance $((V,E),\Sigma,\Phi)$ to a node-weighted graph $(V^{\prime},E^{\prime},w:V^{\prime}\rightarrow\mathbb{Q}^{+})$ as follows.
@@
$E^{\prime}=\{((u,A_{1}),(v,A_{2})\mid(A_{1}\times A_{2})\cap\Phi(e)=\emptyset\text{ for an edge }e\text{ on }u,v\}.$
We can understand this as giving us a graph whose vertices are nodes of the game, labelled by a set of possible colourings. The edges ensure that two vertices cannot be in an independent set together if no possbile colourings of the game nodes from the associated set satisfies the corresponding constraint. As shown in [26, Chapter 5], this is a reduction from $\mathrm{Gap}\textsc{Irreg2to2}_{j,q}(1-\epsilon^{\prime},\delta^{\prime})$ to $\mathrm{GapIS}(1-\frac{1}{\sqrt{2}}-\delta,\epsilon)$.
@@
We now show that this reduction can be defined as a first-order interpretation. The dimension of the reduction is $d=3+|\Sigma|+Q^{|\Sigma|}$. A tuple $t\in V^{d}$ is to be understood as coding one instance of the pair $(v,A)$. As before, the first two elements of $t$ are used to code bits 0 and $1$. The third component is the vertex $v$. The next $|\Sigma|$ components are bits coding the set $A$, and the final $Q^{|\Sigma|}$ elements are a binary identifier giving a value up to $W(v,A)$, thus ensuring that we have $W(v,A)$ instances of the vertex $(v,A)$.
@@
To define the edge relation, consider any 22-to-22 relation $\Pi\subseteq\Sigma\times\Sigma$. We construct a first-order formula $\Pi(x,y)$, where $x$ and $y$ are $d$-tuples, that says that if $x$ and $y$ code vertices $(u,A)$ and $(v,B)$ then there are elements $s\in A$ and $t\in B$ such that $(s,t)\in\Pi$.
@@
$\phi_{E}(x,y):=\exists zC(z)\land\bigvee_{\Pi}[\Phi_{\Pi}(x_{3},y_{3},z)\land\Pi(x,y)].$
In other words, there exists a constraint $z$, in the general 2-to-2 games universe, that is a $\Pi$-constraint between the underlying 2-to-2 nodes $x_{3},y_{3}$ for some 2-to-2 constraint $\Pi$, with the attached set of labels $(A_{x},A_{y})$ satisfying the set-property $(A_{x}\times A_{y})\cap\Phi(e)=\emptyset$.